# Cosmic superstrings, metastable strings and ultralight primordial black holes: from NANOGrav to LIGO and beyond

### v1

@@ III Imprints of ultralight PBHs on GWs from cosmic (super/metastable) strings
$\dot{M}_{\rm BH}=\frac{\mathcal{G}g_{*B}(T_{\rm BH})}{30720\pi}\frac{M_{Pl}^{4}}{M_{\rm BH}^{2}},$ (III.3)
where $M_{Pl}=1.22\times 10^{19}$ GeV, $\mathcal{G}\simeq 3.8$ is a grey body factor and $g_{*B}(T_{\rm BH})\simeq 100$ being the number of relativistic degrees of freedom below the PBH temperature $T_{\rm BH}$ with Standard Model (SM) containing left-handed neutrinos [98]. The quantity $\tilde{H}$ and the scale factor $a$ evolve as
@@
$T_{\rm ev}=\left(\frac{5M_{Pl}^{2}}{\pi^{3}g_{*}(T_{\rm ev})\tau^{2}}\right)^{1/4},$ (III.5)
@@
$\Omega_{\rm GW}(f)\simeq\Omega_{\rm GW}^{\rm plt,\infty}\left(\frac{f}{f_{\rm brk}^{\rm PBH}}\right)^{-1/3}\Theta\left[f-f_{\rm brk}^{\rm PBH}\right],$ (III.8)
@@
$\Omega_{\rm GW}^{\rm plt,\infty}\simeq 3.1\times 10^{-8}\,\Upsilon^{-1}\left(\frac{G\mu}{10^{-7}}\right)^{1/2},\,\,\begin{cases}\Upsilon=P\,\,{\rm for\,\,superstrings}\\ \Upsilon=1\,\,{\rm for\,\,metastable\,strings.}\end{cases}$ (III.9)
Using Eq.(III.7), Eq.(III.8) can be re-expressed as
$\Omega_{\rm GW}(f)\simeq\Omega_{\rm GW}^{\rm plt,\infty}\Omega_{\rm GW}^{\rm mod,PBH},$ (III.10)
@@
$\Omega_{\rm GW}^{\rm mod,PBH}\simeq 0.1\,\left(\frac{10^{-12}}{G\mu}\right)^{1/6}\left(\frac{M_{\rm BH}}{4.1\times 10^{7}\,g}\right)^{-1/2}\left(\frac{f}{25\rm Hz}\right)^{-1/3}$ (III.11)
captures the modulation in the GW spectra because of PBH domination. The absence of PBHs corresponds to $\Omega_{\rm GW}^{\rm mod,PBH}=1$. In which case, Eq.(III.9) and the non-observation of SGWB by LIGO-O3 run; i.e., $\Omega_{\rm GW}^{\rm any\,model}\lesssim 1.7\times 10^{-8}$ at $f_{\rm LIGO}=25$ Hz, implies the following:
${\rm without\,the\,PBH,\,LIGO-O3\,excludes\,\,}\begin{cases}P\lesssim 1.8\,\left(\frac{G\mu}{10^{-7}}\right)^{1/2}\,\,{\rm for\,\,superstrings,}\\ G\mu\gtrsim 3\times 10^{-8}\,\,{\rm for\,\,metastable\,strings.}\end{cases}$ (III.12)
Figure 2: When PBHs dominate in the early universe: Left (right): Gravitational wave spectra for three benchmark points [see Table 2, 2] from each category discussed in the text are shown for superstrings (metastable strings). We consider GWs radiated from loops only for metastable strings, discarding contribution from string segments. Each benchmark GW spectrum is consistent with PTA data and evades LIGO-O3 bound on stochastic gravitational waves.
Though Eq.(III.12) is the most interesting region to explore with non-standard cosmological evolution, e.g., with PBHs, for a comprehensive discussion, we present our numerical results by dividing parameter space into three different categories (all cases are consistent with PTA data):
@@
Here $\Omega_{\rm GW}^{\rm LVK-D}=2\times 10^{-9}$ at $f=25$ Hz is the maximum projected sensitivity (minimum testable value of $\Omega_{\rm GW}$ ) reach of the LVK-D. Note that Category A corresponds to recovering the excluded parameter space in Eq.(III.12), for which one must require a non-standard cosmological evolution, e.g., by PBH. For Category B, PBHs are not essential because even the scale-invariant amplitude does not contradict the LIGO-O3 bound. A PBH domination however can be tested for $f_{\rm brk}^{\rm PBH}\lesssim 25\,{\rm Hz}$ and $\Omega_{\rm GW}^{\rm LIGO-O3}\gtrsim\Omega_{\rm GW}^{\rm plt,\infty}\Omega_{\rm GW}^{\rm mod,PBH}(f=25\,{\rm Hz})\gtrsim\Omega_{\rm GW}^{\rm LVK-D}$. For Category C, though $\Omega_{\rm GW}$ is consistent with the PTA data, the LVK-D cannot test the spectrum because the amplitude is smaller than its sensitivity reach. However, detectors like the Einstein Telescope (ET) can potentially test such signals. We quantify this class of signals by computing Signal-to-Noise-Ratio at ET. In Fig.2, we show the benchmark GW spectrum (summing over larger modes) from each category belonging to both string classes. The benchmark parameters are tabulated in Table 2 (Superstrings) and Table 2 (Metastable strings). In what follows, we present detailed numerical results including a more rigorous parameter space scan to find an approximate mass range of PBHs that can produce such signals.
@@ III.1 Numerical results: PBH mass ranges for PTA GW signals evading LIGO-O3 bound
Figure 3: All plots are for superstrings. Top panel, left: The coloured regions marked with different PBH mass $M_{\rm BH}$ on the $G\mu-P$ plane, correspond to the GW spectra with plateau amplitude $\Omega_{\rm GW}^{\rm plt,\infty}\gtrsim\Omega_{\rm GW}^{\rm LIGO-O3}\equiv 1.7\times 10^{-8}$ and testable with LVK-D at 25 Hz. In the absence of PBHs, the regions would be excluded by LIGO-O3 because in that case, the GW spectra do not fall at $f\ll 25$ Hz, and the scale-invariant amplitudes contradict the LIGO bound. The regions inside the gray contours are consistent with the NANOGrav 15yrs data. Middle: The colored regions correspond to the GW spectra $\Omega_{\rm GW}^{\rm LIGO-O3}\gtrsim\Omega_{\rm GW}^{\rm plt,\infty}\gtrsim\Omega_{\rm GW}^{\rm LVK-D}\equiv 2\times 10^{-9}$. Even in the absence of PBHs, these regions do not contradict the LIGO-O3 bound. But in that case, LVK-D would probe a scale-invariant spectrum. Right: These regions correspond to GWs with smaller amplitude at the LIGO frequency band so that LVK-D can not test them, but detectors such as ET can. For the chosen benchmark point from each category (Table. 2) the GW spectrum is shown in Fig.2 (left). Bottom panel: Zoomed in GW spectra in the PTA band against the NANOGrav posteriors (the violins) and in the LIGO frequency band.
Figure 4: All plots are for metastable strings. Top panel, left: The coloured regions marked with different PBH mass $M_{\rm BH}$ on the $G\mu-\kappa$ plane, correspond to the GW spectra with plateau amplitude $\Omega_{\rm GW}^{\rm plt,\infty}\gtrsim\Omega_{\rm GW}^{\rm LIGO-O3}\equiv 1.7\times 10^{-8}$ and testable with LVK-D at 25 Hz. In the absence of PBHs, the regions would be excluded by LIGO-O3 because in that case, the GW spectra do not fall at $f\ll 25$ Hz, and the scale-invariant amplitudes contradict the LIGO bound. The regions inside the gray contours are consistent with the NANOGrav 15yrs data. Middle: The colored regions correspond to the GW spectra $\Omega_{\rm GW}^{\rm LIGO-O3}\gtrsim\Omega_{\rm GW}^{\rm plt,\infty}\gtrsim\Omega_{\rm GW}^{\rm LVK-D}\equiv 2\times 10^{-9}$. Even in the absence of PBHs, these regions do not contradict the LIGO-O3 bound. Right: These regions correspond to GWs with smaller amplitude at the LIGO frequency band so that LVK-D can not test them, but detectors such as ET can. For the chosen benchmark point from each category (Table 2) the GW spectrum is shown in Fig.2 (right). Bottom panel: Zoomed in GW spectra in the PTA band against the NANOGrav posteriors (the violins) and in the LIGO frequency band.
@@ III.2 Numerical results: Expected Signal-to-Noise-Ratio at the Einstein Telescope
The standard way to assess the detectability of predicted GW signal in the future planned detectors is to compute the Signal-to-Noise-Ratio (SNR) [99, 100]. Because the PTA GWs are of strong amplitudes, a generic expectation from any such signal spanning a wide frequency range would be to obtain a large SNR at the high-frequency detectors. In the present case, because PBH domination reduces the signal amplitude at higher frequencies, it is useful to quantify the detectability of such signals with SNR. Note that, in our case, the higher the frequency smaller the amplitude of GWs. One gets the most reduced signal at higher frequencies for Category C. We therefore assess those signals by computing SNR at the Einstein Telescope. The SNR is defined as [99, 100]
@@
where $\Omega_{\rm noise}(f)$ is the noise spectrum of the detector and $f_{\rm min}(f_{\rm max})$ are minimum (maximum) accessible frequency. In Fig.5, we present the SNR@ET for PTA GWs radiated from superstrings (metastable strings) in the upper (lower) panel for the benchmark values $M_{\rm BH}=10^{7,8}$ g. Note that even though at the ET band the signal strength reduces, the predicted SNR is pretty high. This is owing to the required strong amplitude GWs at the PTA band plus the maximally allowed value of PBH mass which is around $M_{\rm BH}\sim 10^{8}\,g-$the BBN constraint.

### v2

@@ III Imprints of ultralight PBHs on GWs from cosmic (super/metastable) strings
$\dot{M}_{\rm BH}=\frac{\mathcal{G}g_{*B}(T_{\rm BH})}{30720\pi}\frac{M_{\rm Pl}^{4}}{M_{\rm BH}^{2}},$ (III.3)
where $M_{\rm Pl}=1.22\times 10^{19}$ GeV, $\mathcal{G}\simeq 3.8$ is a grey body factor and $g_{*B}(T_{\rm BH})\simeq 100$ being the number of relativistic degrees of freedom below the PBH temperature $T_{\rm BH}$ with Standard Model (SM) containing left-handed neutrinos [104]. The quantity $\tilde{H}$ and the scale factor $a$ evolve as
@@
$T_{\rm ev}=\left(\frac{5M_{\rm Pl}^{2}}{\pi^{3}g_{*}(T_{\rm ev})\tau^{2}}\right)^{1/4},$ (III.5)
@@
Note that the expression for $f_{\rm brk}^{\rm PBH}$ is strictly valid under the assumption of an instantaneous growth of $A_{\beta}$ (from 0.39 to 5.4) as the string network transitions from the PBH-dominated epoch to the radiation-dominated era and the loop number density given in Eq.(II.3) which is a special case of the generic expression of $n_{\omega}(t,l_{k})$ given in Eq.(43) of Ref.[47] under the assumption that the equation of state does not change in time. In reality, either of these is not true during the transition period and must be appropriately considered in the numerical computation.
The correct loop number density and the gradual growth of $A_{\beta}$ (this parameter is defined as $C(t)$ in Ref.[47] in Eq.(45)) approaching its instantaneous value can be tracked accurately using Eq.(III.1), Eq.(III.2), and Eq.(III.4) (note that this equation tracks the time evolution of the scale factor to be fed in Eq.(43) of Ref.[47]) and the velocity-dependent one-scale (VOS) equations of motion [81, 82, 83, 84]:
$\frac{dL}{dt}=HL(1+\tilde{v})^{2}+\frac{1}{2}\tilde{c}\tilde{v},$ (III.8)
$\frac{d\tilde{v}}{dt}=1-\tilde{v}^{2}\left(\frac{k(\tilde{v})}{L}-2H\tilde{v}\right),$ (III.9)
where $\tilde{v}$ is the root-mean-squared (RMS) velocity of the long strings, $c=0.23\pm 0.04$ and the function $k(\tilde{v})$ which phenomenologically accounts for the effects of small structures on the long strings is given by
$k(\tilde{v})=\frac{2\sqrt{2}}{\pi}(1-\tilde{v}^{2})(1+2\sqrt{2}\tilde{v}^{3})\frac{1-8\tilde{v}^{6}}{1+8\tilde{v}^{6}}.$ (III.10)
Parametrising $L=\xi(t)t$ one defines $A_{\beta}$ as $A_{\beta}=\frac{\tilde{c}\tilde{v}}{\sqrt{2}\xi^{3}}$, where $\xi(t)=\alpha(t)/\alpha_{L}$ with $\alpha_{L}\simeq 0.37$. So far we have assumed that Eq.(III.8) and Eq.(III.9) reach their attractor solutions ($\tilde{v}$ and $\xi$ are constant) instantly during the change of cosmology. In our case, precise evolution of $A_{\beta}(z)$ can be obtained upon recasting Eq.(III.8) and Eq.(III.9) as
$\frac{dt}{dz}$ $=$ $\frac{1}{\tilde{H}z},$ (III.11)
$\frac{d\xi}{dz}$ $=$ $\frac{1}{t\tilde{H}z}\left(Ht\xi(1+\tilde{v})^{2}+\frac{1}{2}\tilde{c}\tilde{v}-\xi\right),$ (III.12)
$\frac{d\tilde{v}}{dz}$ $=$ $\frac{1-\tilde{v}^{2}}{\tilde{H}z}\left(\frac{k(\tilde{v})}{\xi t}-2H\tilde{v}\right)$ (III.13)
and solving them together with Eq.(III.1) and Eq.(III.2).
Figure 2: Evolution of normalized energy densities and $A_{\beta}$ with $z$ in presence of ultralight PBH with initial mass $M_{\rm BH}=4.1\times 10^{7}$ g. The blue and the red lines representing $A_{\beta}$ correspond to two different initial conditions.
In Fig. 2, we present the evolution of $A_{\beta}$ as a function of $z$ for two distinct initial conditions: $\tilde{v}_{\rm in}=0.1$ and $\xi_{\rm in}=0.1,1$, represented by the red and blue curves, respectively. The figure clearly illustrates that, in both scenarios, $A_{\beta}$ eventually converges to the attractor solution. However, the transitions are gradual in both cases: the shift from the initial radiation epoch to the PBH-dominated epoch occurs around $z\sim 10^{5}$, followed by a transition from the PBH-dominated epoch to the latest radiation era near $z\sim 10^{14}$. The later transition is important in our discussion.
After appropriately taking into account the loop number density (Eq.(43) of Ref.[47]) and variation of $A_{\beta}$ presented in Fig.2, we find that an approximate analytical relation for the spectral break frequency can be written as
$f_{\rm brk}^{\rm PBH}\simeq 0.008\,{\rm Hz}\left(\frac{10^{-12}}{G\mu}\right)^{1/2}\left(\frac{M_{\rm BH}}{4.1\times 10^{7}\,g}\right)^{-3/2}\,$ (III.14)
accounting for a percent level deviation from the near-scale-invariant plateau. For further computations, we will use the spectral break frequency defined in Eq. (III.14), which is smaller by order of magnitude compared to the instantaneous case shown in Eq. (III.7).
@@
$\Omega_{\rm GW}(f)\simeq\Omega_{\rm GW}^{\rm plt,\infty}\left(\frac{f}{f_{\rm brk}^{\rm PBH}}\right)^{-1/3}\Theta\left[f-f_{\rm brk}^{\rm PBH}\right],$ (III.15)
@@
$\Omega_{\rm GW}^{\rm plt,\infty}\simeq 3.1\times 10^{-8}\,\Upsilon^{-1}\left(\frac{G\mu}{10^{-7}}\right)^{1/2},\,\,\begin{cases}\Upsilon=P\,\,{\rm for\,\,superstrings}\\ \Upsilon=1\,\,{\rm for\,\,metastable\,strings.}\end{cases}$ (III.16)
Using Eq.(III.7), Eq.(III.15) can be re-expressed as
$\Omega_{\rm GW}(f)\simeq\Omega_{\rm GW}^{\rm plt,\infty}\Omega_{\rm GW}^{\rm mod,PBH},$ (III.17)
@@
$\Omega_{\rm GW}^{\rm mod,PBH}\simeq 0.068\,\left(\frac{10^{-12}}{G\mu}\right)^{1/6}\left(\frac{M_{\rm BH}}{4.1\times 10^{7}\,g}\right)^{-1/2}\left(\frac{f}{25\,\rm Hz}\right)^{-1/3}$ (III.18)
captures the modulation in the GW spectra because of PBH domination. The absence of PBHs corresponds to $\Omega_{\rm GW}^{\rm mod,PBH}=1$. In which case, Eq.(III.16) and the non-observation of SGWB by LIGO-O3 run; i.e., $\Omega_{\rm GW}^{\rm any\,model}\lesssim 1.7\times 10^{-8}$ at $f_{\rm LIGO}=25$ Hz, implies the following:
${\rm without\,the\,PBH,\,LIGO-O3\,excludes\,\,}\begin{cases}P\lesssim 1.8\,\left(\frac{G\mu}{10^{-7}}\right)^{1/2}\,\,{\rm for\,\,superstrings,}\\ G\mu\gtrsim 3\times 10^{-8}\,\,{\rm for\,\,metastable\,strings.}\end{cases}$ (III.19)
Figure 3: When PBHs dominate in the early universe: Left (right): Gravitational wave spectra for three benchmark points [see Table 2, 2] from each category discussed in the text are shown for superstrings (metastable strings). We consider GWs radiated from loops only for metastable strings, discarding contribution from string segments. Each benchmark GW spectrum is consistent with PTA data and evades LIGO-O3 bound on stochastic gravitational waves. To produce these GW spectra and all the other spectra in subsequent figures we sum $10^{3}$ modes and the extrapolate it to higher frequencies keeping $f^{-1/3}$ falling.
Though Eq.(III.19) is the most interesting region to explore with non-standard cosmological evolution, e.g., with PBHs, for a comprehensive discussion, we present our numerical results by dividing parameter space into three different categories (all cases are consistent with PTA data):
@@
Here $\Omega_{\rm GW}^{\rm LVK-D}=2\times 10^{-9}$ at $f=25$ Hz is the maximum projected sensitivity (minimum testable value of $\Omega_{\rm GW}$ ) reach of the LVK-D. Note that Category A corresponds to recovering the excluded parameter space in Eq.(III.19), for which one must require a non-standard cosmological evolution, e.g., by PBH. For Category B, PBHs are not essential because even the scale-invariant amplitude does not contradict the LIGO-O3 bound. A PBH domination however can be tested for $f_{\rm brk}^{\rm PBH}\lesssim 25\,{\rm Hz}$ and $\Omega_{\rm GW}^{\rm LIGO-O3}\gtrsim\Omega_{\rm GW}^{\rm plt,\infty}\Omega_{\rm GW}^{\rm mod,PBH}(f=25\,{\rm Hz})\gtrsim\Omega_{\rm GW}^{\rm LVK-D}$. For Category C, though $\Omega_{\rm GW}$ is consistent with the PTA data, the LVK-D cannot test the spectrum because the amplitude is smaller than its sensitivity reach. However, detectors like the Einstein Telescope (ET) can potentially test such signals. We quantify this class of signals by computing Signal-to-Noise-Ratio at ET. In Fig.3, we show the benchmark GW spectrum (summing over larger modes) from each category belonging to both string classes. The benchmark parameters are tabulated in Table 2 (Superstrings) and Table 2 (Metastable strings). In what follows, we present detailed numerical results including a more rigorous parameter space scan to find an approximate mass range of PBHs that can produce such signals.
@@ III.1 Numerical results: PBH mass ranges for PTA GW signals evading LIGO-O3 bound
Figure 4: All plots are for superstrings. Top panel, left: The coloured regions marked with different PBH mass $M_{\rm BH}$ on the $G\mu-P$ plane, correspond to the GW spectra with plateau amplitude $\Omega_{\rm GW}^{\rm plt,\infty}\gtrsim\Omega_{\rm GW}^{\rm LIGO-O3}\equiv 1.7\times 10^{-8}$ and testable with LVK-D at 25 Hz. In the absence of PBHs, the regions would be excluded by LIGO-O3 because in that case, the GW spectra do not fall at $f\ll 25$ Hz, and the scale-invariant amplitudes contradict the LIGO bound. The regions inside the gray contours are consistent with the NANOGrav 15yrs data. Middle: The colored regions correspond to the GW spectra $\Omega_{\rm GW}^{\rm LIGO-O3}\gtrsim\Omega_{\rm GW}^{\rm plt,\infty}\gtrsim\Omega_{\rm GW}^{\rm LVK-D}\equiv 2\times 10^{-9}$. Even in the absence of PBHs, these regions do not contradict the LIGO-O3 bound. But in that case, LVK-D would probe a scale-invariant spectrum. Right: These regions correspond to GWs with smaller amplitudes at the LIGO frequency band, so LVK-D can not test them, but detectors such as ET can. For the chosen benchmark point from each category (Table. 2) the GW spectrum is shown in Fig.3 (left). Bottom panel: Zoomed in GW spectra in the PTA band against the NANOGrav posteriors (the violins) and in the LIGO frequency band.
Figure 5: All plots are for metastable strings. Top panel, left: The coloured regions marked with different PBH mass $M_{\rm BH}$ on the $G\mu-\kappa$ plane, correspond to the GW spectra with plateau amplitude $\Omega_{\rm GW}^{\rm plt,\infty}\gtrsim\Omega_{\rm GW}^{\rm LIGO-O3}\equiv 1.7\times 10^{-8}$ and testable with LVK-D at 25 Hz. In the absence of PBHs, the regions would be excluded by LIGO-O3 because in that case, the GW spectra do not fall at $f\ll 25$ Hz, and the scale-invariant amplitudes contradict the LIGO bound. The regions inside the gray contours are consistent with the NANOGrav 15yrs data. Middle: The colored regions correspond to the GW spectra $\Omega_{\rm GW}^{\rm LIGO-O3}\gtrsim\Omega_{\rm GW}^{\rm plt,\infty}\gtrsim\Omega_{\rm GW}^{\rm LVK-D}\equiv 2\times 10^{-9}$. Even in the absence of PBHs, these regions do not contradict the LIGO-O3 bound. Right: These regions correspond to GWs with smaller amplitudes at the LIGO frequency band, so LVK-D can not test them, but detectors such as ET can. For the chosen benchmark point from each category (Table 2) the GW spectrum is shown in Fig.3 (right). Bottom panel: Zoomed in GW spectra in the PTA band against the NANOGrav posteriors (the violins) and in the LIGO frequency band.
@@ III.2 Numerical results: Expected Signal-to-Noise-Ratio at the Einstein Telescope
The standard way to assess the detectability of predicted GW signal in the future planned detectors is to compute the Signal-to-Noise-Ratio (SNR) [105, 106, 107, 108, 109, 110]. Because the PTA GWs are of strong amplitudes, a generic expectation from any such signal spanning a wide frequency range would be to obtain a large SNR at the high-frequency detectors. In the present case, because PBH domination reduces the signal amplitude at higher frequencies, it is useful to quantify the detectability of such signals with SNR. Note that, in our case, the higher the frequency, the smaller the amplitude of GWs. One gets the most reduced signal at higher frequencies for Category C. We assess those signals by computing SNR at the Einstein Telescope (ET). We quantify the SNR using the most general (for arbitrary signal strength) optimal filter function [110]
$\tilde{Q}(f)=\frac{U(f)V(f)^{*}-U(f)^{*}W(f)}{|V(f)|^{2}-W(f)^{2}}$ (III.20)
that returns SNR as [110]
${\rm SNR}=\sqrt{2T_{\rm obs}}\left[\int_{-\infty}^{+\infty}\frac{df}{2}\frac{|U(f)|^{2}}{|U(f)|^{2}+W(f)}\right]^{1/2},$ (III.21)
where the frequency-dependent functions are given by
$U(f)$ $=$ $C_{IJ}(f)+\delta_{IJ}N_{I}(f),$ (III.22)
$V(f)$ $=$ $C_{IJ}^{2}+\delta_{IJ}N_{I}\big{(}N_{I}+2C_{II}\big{)},$ (III.23)
$W(f)$ $=$ $C_{II}C_{JJ}+C_{II}N_{J}+C_{JJ}N_{I}+N_{I}N_{J}$ (III.24)
with
$C_{IJ}(f;t,t)$ $=$ $\int\frac{d\Omega}{4\pi}S_{h}(f)F_{IJ}(f,\Omega;t,t),$ (III.25)
$\braket{\tilde{n}^{*}_{I}(f)\tilde{n}_{J}(f^{\prime})}$ $=$ $\frac{1}{2}\delta_{IJ}\delta(f-f^{\prime})N_{I}(f)$ (III.26)
$\langle\tilde{h}^{*}_{A}(f,\Omega)\tilde{h}_{A^{\prime}}(f^{\prime},\Omega^{\prime})\rangle$ $=$ $\frac{1}{2}\delta(f-f^{\prime})\delta^{2}(\Omega,\Omega^{\prime})\frac{\delta_{AA^{\prime}}}{4\pi}S_{h}(|f|,\Omega).$ (III.27)
The indices $I,J$ represent the different detectors used to detect the signal, and $F_{IJ}$ is the antenna pattern function, which characterizes the overlap in the detectors’ responses to the signal (related to the overlap reduction function [109]). The term $\tilde{n}$ denotes the noise, while $\tilde{h}$ represents the GW signal. The indices $A,A^{\prime}$ indicate the polarization states of the GW signal, and $\Omega$ denotes the direction from which the GW originates. The strain power spectral density $S_{h}(f)$ is related to the $\Omega_{\text{GW}}(f)$ as
$\Omega_{\text{GW}}(f)=\frac{4\pi}{3}\frac{f^{3}S_{h}(f)}{H_{0}^{2}}.$ (III.28)
Defining $C_{IJ}\equiv S_{h}(f)\Gamma_{IJ}$, for ET detector network SNR in Eq.(III.21) simplifies to
@@
where we have assumed equal noise for all the detectors and defined $S_{n}(f)=\frac{N}{\sqrt{3}\Gamma_{\rm ET}}$ [106], which is related to $\Omega_{\rm noise}$ following Eq.(III.28). Note that the SNR defined in Eq.(III.29), reduces to the standard formula [105, 106] used vastly in the literature in the weak signal limit ($\Omega_{\rm GW}\ll\Omega_{\rm noise}$). In Fig.6, we present the ${\rm SNR}_{\rm ET}$ for PTA GWs radiated from superstrings (metastable strings) in the upper (lower) panel for the benchmark values $M_{\rm BH}=10^{7,8}$ g. Note that even though at the ET band the signal strength reduces, the predicted SNR is pretty high. This is owing to the required strong amplitude GWs at the PTA band plus the maximum value of PBH mass, around $M_{\rm BH}\sim 10^{8}\,g-$the BBN constraint.